OpenStudy (liddoislander):

Use the distributive property to simplify the following expression: y(y + 13)= A.2y + 13 B.y2 + 13 C.y2 + 13y D.2y2 + 13

OpenStudy (bibby):

the distributive property would look like this a(b+c)= ab + ac. You distribute the a to each term. as such y(y+13) = y*y + 13*y. which answer does that look like>

OpenStudy (anonymous):

\[y ^{2}+13y\]

OpenStudy (bibby):

Thank you for just giving the answer and avoiding an explanation. You really help the community

OpenStudy (liddoislander):

thank you ! you guys are amazing ! & thank you bibby for explaining lol

OpenStudy (anonymous):

How about you keep your thoughts to yourself, and give her your explanation to help the community rather than be sassy. It helps society a lot.

OpenStudy (bibby):

I'm not being sassy, You're doing his/her homework instead of teaching them the math they need to do their own homework and pass their tests. Go to yahoo answers if you want to just spoonfeed some kids homework for them